OpenStudy (anonymous):

Suppose you roll a six-sided die two times hoping to get two numbers whose sum is even. What is the sample space? How many favorable outcomes are there?

OpenStudy (anonymous):

9/36

OpenStudy (anonymous):

S = 36

OpenStudy (anonymous):

sample space has 36 elements !!! 11 12 13 14 15 16 21 --------------- 31 ---------------- 41 ---------------- 51 ---------------- 61 ---------------- make it yourself and then see how many of them fulfil your condition and then get your answer

OpenStudy (anonymous):

sample space is 36 favorable outcomes are 12/36 = 1/3

OpenStudy (anonymous):

favourable outcomes are 18 !!!

OpenStudy (anonymous):

well it depends on whether you allow for repetition or not

OpenStudy (anonymous):

27/36

OpenStudy (anonymous):

@kanwar 27 favourable outcomes ????? pls count again

OpenStudy (anonymous):

it says nothing about repeating all it says is sum has to be even

OpenStudy (anonymous):

favorable out comes are 1-1 1-3 1-5 2-2 2-4 2-6 3-3 3-5 4-4 4-6 5-5 6-6

OpenStudy (anonymous):

11 13 15 22 24 26 31 33 35 42 44 46 51 53 55 62 64 and 66 i.e 18 fAVOURABLE OUTCOMES

OpenStudy (anonymous):

@dhatraditya u missed some !!!

OpenStudy (anonymous):

Yes, harkirat is correct, you have to count the inverse occurences (ie 1-3 and 3-1) since they are separate events.

OpenStudy (anonymous):

right, harkirat is right, but I didn't consider the repetitious ones. I am assuming 1-3 is the same as 3-1

OpenStudy (anonymous):

do a sample space diagram. numbers on 1st dice. 1 2 3 4 5 6 ________________________________________________________________ n 1| 1+1=2 2+1=3 4 5 6 7 u 2 | m o n 2| 3 4 5 6 7 8 b n d | e 3| 4 5 6 7 8 9 r d | s i 4| 5 6 7 8 9 10 c | e 5| 6 7 8 9 10 11 | 6| 7 8 9 10 11 12 so we are adding the top number to the side number to get the entries in the table. Total amount of even numbers in the table is 18 total amount of entries is in the table is 36. Probability of getting an even answer on summing the die totals =18/36 = 1/2

OpenStudy (anonymous):

if not, and you treat the two dice as distinct events, then yes, harkirat is right

OpenStudy (anonymous):

no no 13 is never same as 31, they r distinctly separate events

OpenStudy (anonymous):

Also it makes intuitive sense that half of the possible outcomes are even when both dice have equal numbers even and odd numbers on them.